OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (-8, 0), (1, 5)

OpenStudy (anonymous):

m=(5-0)/(1--8)=5/9 y=(5/9)(x--8)+0 =y=(5/9)x+(40/9) =(5/9)x+(40/9)-y clear fractions 5x+40-9y